# An ultra-broadband photonic-chip-based traveling-wave parametric amplifier

### v1

@@
Figure 3: Amplification of optical frequency combs using a GaP optical TWPA. (a) Schematic of amplification and frequency conversion of EO frequency comb (top) and soliton microcomb (bottom). (b) Amplified EO comb spectra (red) and fiber-coupled input spectrum of OPA (blue) showing more than 20 dB fiber-to-fiber net gain. Envelopes are added for better visibility. (c) Same as (b) but for amplification of 100 GHz microcomb produced by single soliton state in a Si3N4 microring cavity. The inset shows the appearance of frequency-converted idler comb lines as a result of DFWM. Increased noise floor at longer wavelengths on both spectra is due to the proximity to the end of the OSA measurement range. The resolution bandwidth of the OSA is 0.02 nm in both experiments.
To highlight the application potential of such a broadband optical TWPA, we perform additional amplification experiments using low-power and high-repetition-rate frequency comb sources as signals. In Fig. 3(a), we schematically show amplification and idler generation for two types of injected frequency combs, a narrowband electro-optic (EO) comb [38] and a broadband dissipative Kerr soliton comb [41]. The EO comb is centered around 1600 nm and has a 16 GHz line spacing and a total signal power of 10.5 $\mu$W. When transmitted through the amplifier together with a 4.12 W pump at 1550 nm, we observe more than 20 dB fiber-to-fiber net gain (Fig. 3(b)). The reference dataset (blue) is measured by bypassing the lensed fibers and the photonic chip. Note that, due to the nearly instantaneous nature of parametric amplification, we do not use a dispersive fiber compression stage during the EO comb preparation before the amplifier to avoid high signal peak powers and amplifier saturation. We perform the same experiment with the frequency comb formed by a dissipative Kerr soliton state [41] generated in a Si3N4 microring resonator. The total power of the soliton comb is 76 $\mu$W, and the repetition rate is 100 GHz. In the spectral region of the highest gain, far from the soliton pump located at 1552.8 nm, the input power of individual lines decreases to the level of a few nW, while lines with the highest initial power, located closer to the pump, experience smaller gain. As expected, an idler comb mirrored with respect to the pump frequency is generated, and the line density of the output spectrum is doubled (Fig. 3(c), red). Additional lines can be observed throughout the entire spectrum and, most notably, closer to the pump; we attribute their appearance to numerous FWM processes between various interacting lines. Additional lines that are located at frequencies far outside the bandwidth of the input comb arise due to cascaded non-degenerate FWM processes. A small feature in the amplified spectra around 1651 nm is observed with all available waveguides and can be attributed to Raman scattering in GaP [37, 42]. The flattened spectra resulting from amplification provide clear evidence that the GaP OPA can handle the simultaneous input of a few hundred lines over a broad bandwidth, and the parametric gain remains as high as for a single input frequency if operated in the small-signal regime. Moreover, we again observe amplification of very small input signals, with the lowest comb lines being as weak as 1 nW, something that is not possible with typical EDFAs. The amplification of such small signals may be especially attractive for applications such as LiDAR [43] or free-space communications.
@@
Figure 4: Coherent data transmission experiment. (a) Simplified schematic of the experimental setup used in the communication experiment AWG: arbitrary waveform generator. FBG: fiber Bragg grating. WDM: wavelength-division multiplexer. OSC: oscilloscope. (b) Calibrated optical spectra measured at the output of the chip using the same OSA as described above (not shown here in (a)); resolution bandwidth is set to 0.2 nm for a faster spectra acquisition. (c, d, e) Constellation diagrams and modulation error ratios (MERs) for three measurements, respectively: fiber-to-fiber reference, amplified signal, generated idler.
@@
To further demonstrate the potential of the GaP OPA for use in coherent communication, we set up a 10 GBd communication line, as shown in Fig. 4(a). We send a QPSK-encoded pseudo-random bit sequence at a signal wavelength of 1605 nm (L-band) to our OPA and then perform heterodyne measurements of the amplified signal. Additionally, using another laser as a local oscillator, we measure modulation transfer to the idler located at approximately 1502 nm (S-band). The signal wavelength is chosen to be in the region of maximum gain, while also considering the limitations of the available equipment. Fig. 4(b) shows corrected OPA output spectra with the pump switched on and off. The internal gain achieved exceeds 30 dB, which is enough to compensate for coupling and propagation losses on the chip (which barely exceed 10 dB in total at the selected signal wavelength) and even for 10 dB of losses on the splitter before the chip and 4 dB of losses in the filtering section and wavelength-division multiplexer, so we achieve a positive net gain accounting for all losses present in the system (see Supplementary Material). The received signal power is 26 $\mu$W and the idler power is 32 $\mu$W; the idler power is higher than the signal power due to lower losses at the idler wavelength. As a reference, we measure the signal directly sent from the transmitter to the receiver through a fiber, bypassing all the other components without attenuation. We digitally post-process and analyze the data and plot constellation diagrams for all three measurements – for the fiber reference, signal, and idler (Figs. 4(C, D, E), respectively). As a figure of merit, for each measurement we estimate the modulation error ratio (MER) [44], defined as the sum of the squares of the magnitude of the ideal symbol vectors divided by the sum of the squares of the magnitude of the symbol error vectors. The idler conversion efficiency is high enough in the GaP OPA to generate an idler of the same power level as that of the amplified signal. This eliminates any need for additional idler post-amplification, as it is sufficiently strong to be used directly for inter-band signal translation.
Discussion. In summary, using an integrated GaP photonic platform, we achieve both high net gain and broadband operation of an OPA. The fiber-to-fiber net gain exceeds 20 dB, and the combined signal and idler gain bandwidth is 140 nm, significantly greater than the bandwidth of EDFAs. We demonstrate the practical application potential of GaP waveguides for the amplification of frequency combs, optical coherent data streams, and weak optical signals over more than 60 dB of dynamic range. Further reducing the optical propagation losses will greatly lower the required pump power and may even allow direct pumping of the OPA with semiconductor lasers.
@@
Our on-chip GaP OPA solves some of the most pivotal challenges that have prevented the widespread adoption of fiber OPA systems. The dispersion of the waveguides can be designed as needed because it is defined lithographically, and the waveguide’s short length avoids a reduction in bandwidth [45] and reduces sensitivity to fabrication imperfections, in contrast to optical fibers that are hundreds of meters long. The low Brillouin gain of integrated waveguides eliminates the technical complexity of adding additional pump phase modulators [19, 20, 15]. In principle, the pump frequency can be freely chosen over a wide frequency range, as it is not restricted by an excitation spectrum as in doped amplifiers. Also, in contrast to existing EDFA technology, the optical TWPA is unidirectional, making it more robust against parasitic reflections. Lastly, the TWPA may achieve a lower noise figure, which can be reduced below the quantum limit if operated in the phase-sensitive regime [16, 34].
Our results signal the emergence of practical optical TWPAs with large bandwidth, high gain, and small footprint that could eventually be employed in future optical communication systems or other applications. Broadband optical TWPAs may be ideal candidates to drive next-generation fiber systems relying on doubly antiresonant hollow-core fibers that offer much increased spectral bandwidth for optical communications compared to silica fibers. Nonlinear distortions in parametrically amplified fiber links may also be mitigated by alternately propagating the signal and phase-conjugated idler through successive fiber link sections [46]. In general, the ability to amplify weak optical signals over a large dynamic range may be decisive for a variety of applications, including coherent communications, LiDAR and sensing.
@@
Funding information: This work was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 812818 (MICROCOMB), by the Swiss National Science Foundation (SNF) under grant number 192293, and by the Air Force Office of Scientific Research under award number FA9550-19-1-0250.
@@
Disclosures: J.R., N.K. and T.J.K. are inventors on European Patent Application No. 24165690.9 submitted by Swiss Federal Institute of Technology Lausanne (EPFL) that covers “Optical parametric amplifier apparatus and optical signal transmission apparatus, and applications thereof”. Other authors declare no conflicts of interest.
Data availability: All experimental datasets and scripts used to produce the plots in this work will be uploaded to Zenodo repository upon publication of this preprint.

### v2

@@
Figure 3: Amplification of optical frequency combs using a GaP optical TWPA. (a) Schematic of amplification and frequency conversion of a EO frequency comb (top) and a dissipative Kerr soliton frequency comb (bottom). (b) Spectra of the amplified (red) and input (blue) EO frequency comb showing more than 20 dB fiber-to-fiber net gain. Envelopes are added for better visibility. (c) Same as (b) but for amplification of a 100-GHz frequency comb produced by a single soliton state in a Si3N4 microring cavity. The inset shows the appearance of frequency-converted idler-comb lines as a result of DFWM. The increased noise floor at longer wavelengths in both spectra is due to the proximity to the end of the OSA measurement range. The resolution bandwidth of the OSA is 0.02 nm in both experiments.
At 2.01 W pump power, the amplification is linear over the entire range of input powers measured, extending over six orders of magnitude and limited only by the capabilities of our measurement setup; in principle, parametric amplification poses no lower limit on the signal input power. We calculate the amplifier noise figure from the measured signal amplification and the spontaneous parametric fluorescence background (see Supplementary Material) [41]. The off-chip noise figure approaches 6 dB in the small-signal gain regime. Since degradation of the signal-to-noise ratio (SNR) at the input facet is irreversible, we take into account the coupling losses of 2.5 dB at the chip input facet and find the on-chip noise figure of less than 4 dB for a wide range of signal powers below saturation (Fig. 2(g)). The degradation of the noise figure at higher signal input powers is attributed to gain saturation; in other words, the signal gain becomes weaker while the parametric fluorescence is still within the small-signal regime and undergoes a higher amplification.
Amplification of optical frequency combs To highlight the application potential of such a broadband optical TWPA, we perform additional amplification experiments using low-power and high-repetition-rate frequency comb sources as signals. In Fig. 3(a), we schematically show amplification and idler generation for two types of injected frequency combs, a narrowband electro-optic (EO) comb [15] and a broadband dissipative Kerr soliton comb [42]. The EO comb is centered around 1600 nm and has a 16-GHz line spacing and a total signal power of 10.5 $\mu$W.
@@
Figure 4: Coherent data transmission experiment. (a) Simplified schematic of the experimental setup used in the communication experiment. AWG: arbitrary waveform generator; FBG: fiber Bragg grating; WDM: wavelength-division multiplexer; OSC: oscilloscope; LO: local oscillator. (b) Calibrated optical spectra measured at the output of the chip using the same OSA as described above (not shown here in (a)). The resolution bandwidth is set to 0.2 nm for faster acquisition. (c, d, e) Constellation diagram and MER for three measurements: fiber-to-fiber reference, amplified signal, and generated idler.
When transmitted through the amplifier together with a 4.12-W pump at 1550 nm, we observe more than 20 dB fiber-to-fiber net gain (Fig. 3(b)). The reference dataset (blue) is measured by bypassing the lensed fibers and the photonic chip. Note that, due to the nearly instantaneous nature of parametric amplification, we do not use a dispersive fiber compression stage during the EO comb preparation before the amplifier to avoid high signal peak powers and amplifier saturation. We perform the same experiment with the frequency comb formed by a dissipative Kerr soliton state [42] generated in a Si3N4 microring resonator. The total power of the soliton comb is 76 $\mu$W, and the repetition rate is 100 GHz. In the spectral region of the highest gain, far from the soliton pump located at 1552.8 nm, the input power of individual lines decreases to the level of a few nW, while lines with the highest initial power, located closer to the pump, experience smaller gain. As expected, an idler comb mirrored with respect to the pump frequency is generated, and the line density of the output spectrum is doubled (Fig. 3(c), red). Additional lines can be observed throughout the entire spectrum and, most notably, closer to the pump; we attribute their appearance to numerous FWM processes between various interacting lines. Additional lines that are located at frequencies far outside the bandwidth of the input comb arise due to cascaded non-degenerate FWM processes. A small feature in the amplified spectra around 1651 nm is observed with all available waveguides and can be attributed to Raman scattering in GaP [14, 43]. The flattened spectra resulting from amplification provide clear evidence that the GaP OPA can handle the simultaneous input of a few hundred lines over a broad bandwidth, and the parametric gain remains as high as for a single input frequency if operated in the small-signal regime. Moreover, we again observe amplification of very small input signals, with the lowest comb lines being as weak as 1 nW, something that is not possible with typical EDFAs. The amplification of such small signals may be especially attractive for applications such as LiDAR [44] or free-space communications.
@@
Coherent communication and modulation transfer To further demonstrate the potential of the GaP OPA for use in coherent communication, we set up a 10 GBd communication line, as shown in Fig. 4(a). We send a QPSK-encoded pseudo-random bit sequence at a signal wavelength of 1605 nm (L-band) to our OPA and then perform heterodyne measurements of the amplified signal. Additionally, using another laser as a local oscillator (LO), we measure modulation transfer to the idler located at approximately 1502 nm (S-band). The signal wavelength is chosen to be in the region of maximum gain, while also considering the limitations of the available equipment. Fig. 4(b) shows the corrected OPA output spectra with the pump switched on and off. The internal gain achieved exceeds 30 dB, which is enough to compensate for coupling and propagation losses on the chip (which barely exceed 10 dB in total at the selected signal wavelength) and even for 10 dB of losses on the splitter before the chip and 4 dB of losses in the filtering section and wavelength-division multiplexer, so we achieve a positive net gain accounting for all losses present in the system (see Supplementary Material). The received signal power is 26 $\mu$W and the idler power is 32 $\mu$W; the idler power is higher than the signal power due to lower losses at the idler wavelength. The LO power is 29.1 mW for the signal measurement and 27.7 mW for the idler. As a reference, we measure the signal directly sent from the transmitter to the receiver through a fiber, bypassing all the other components without attenuation. We digitally post-process and analyze the data and plot constellation diagrams for all three measurements – for the fiber reference, signal, and idler (Figs. 4(c, d, e), respectively). As a figure of merit, for each measurement we estimate the modulation error ratio (MER) [45], defined as the sum of the squares of the magnitude of the ideal symbol vectors divided by the sum of the squares of the magnitude of the symbol error vectors (see Supplementary Material). The idler conversion efficiency is high enough in the GaP OPA to generate an idler of the same power level as that of the amplified signal. This eliminates any need for additional idler post-amplification, as it is sufficiently strong to be used directly for inter-band signal translation.
Discussion In summary, using an integrated GaP photonic platform, we achieve both high net gain and broadband phase-insensitive operation of an OPA. The maximum fiber-to-fiber net gain reaches 25 dB, and the combined signal and idler 10-dB-gain bandwidth is 140 nm, substantially greater than the bandwidth of both EDFAs and existing CW parametric amplification systems [10, 8, 11]. We demonstrate the potential for practical application of GaP waveguides for the amplification of frequency combs, coherent optical data streams, and optical signals over more than 60 dB of dynamic range, fully covering and extending beyond input powers from -26 dBm to 0 dBm, and output powers from -8 dBm to 6 dBm, which are typical requirements in telecommunication technology [46]. The ability to amplify weak optical signals over a large dynamic range may be decisive for a variety of applications beyond coherent communications, including LiDAR and sensing. Moreover, integrated TWPA can be utilized where amplifiers are needed in custom bands, such as 1650 nm for methane detection or in applications like optical coherence tomography requiring amplification at 1300 nm, where no chip-scale solutions based on rare-earth ions currently exist.
Further reducing the optical propagation losses will greatly lower the required pump power and may even allow direct pumping of the OPA with semiconductor lasers.
@@
Our on-chip GaP OPA solves some of the most pivotal challenges that have prevented the widespread adoption of fiber OPA systems. Viewed more broadly, the results mark an example where the performance of nonlinear photonic integrated circuit-based amplifiers surpasses that of legacy fiber-based systems. The dispersion of the waveguide can be designed as needed because it is defined lithographically, and the waveguide’s short length ensures a broad bandwidth [47] and reduces sensitivity to fabrication imperfections, in contrast to optical fibers that are hundreds of meters long. Compared to other photonic-chip-based TWPAs [11, 12], the GaP TWPA features a waveguide length and overall device footprint that are reduced by orders of magnitude. The low Brillouin gain of integrated waveguides eliminates the technical complexity of adding pump phase modulators [7, 8, 26, 48]. In principle, the pump frequency can be freely chosen over a wide frequency range when designing the TWPA, as it is not restricted by an excitation spectrum as in doped amplifiers. Also, in contrast to existing EDFA technology, the optical TWPA is unidirectional, making it more robust against parasitic reflections, which not only remain unamplified but are actually further suppressed due to the propagation losses. Lastly, the TWPA can in principle achieve a lower noise figure than an EDFA and lower than what we have demonstrated here, as it can be operated as a phase-sensitive amplifier [27, 12], reducing the noise below the quantum limit and simultaneously increasing the peak gain by 6 dB.
Our results signal the emergence of compact, high-performance photonic integrated circuit based optical integrated TWPAs with large bandwidth, high gain and small footprint that have the potential to transition from the laboratory into future optical communication systems. Broadband optical TWPAs may find use in a variety of future system configurations. For example, nonlinear distortions in parametrically amplified fiber links may be mitigated by alternately propagating the signal and phase-conjugated idler through successive fiber link sections [49].
@@
Funding information: This work was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 812818 (MICROCOMB), by the Swiss National Science Foundation (SNF) under grant number 216493 (HEROIC), and by the Air Force Office of Scientific Research under award number FA9550-19-1-0250.
@@
Disclosures: J.R., N.K. and T.J.K. are inventors on European Patent Application No. 24165690.9 submitted by Swiss Federal Institute of Technology Lausanne (EPFL) that covers "Optical parametric amplifier apparatus and optical signal transmission apparatus, and applications thereof". Other authors declare no competing interests.
Data availability: All experimental datasets and scripts used to produce the plots in this work will be uploaded to the Zenodo repository upon publication of this preprint.